Projectile Motion


1. A projectile is launched at ground level with an initial speed of 50 0 m/s at an angle of 30.0
1
Expert's answer
2021-09-26T18:49:00-0400

All quantities that can be asked to found in such tasks can be calculated using these formulas:

https://courses.lumenlearning.com/boundless-physics/chapter/projectile-motion/


Let "u = 50m\/s" be the initial speed and "\theta =30\degree" the launch angle. Then, for example, the range of the motion is given as follows:


"R = \dfrac{u^2\cdot \sin2\theta }{g} = \dfrac{50^2\cdot \sin(2\cdot 30\degree)}{9.81} \approx 221 m,"

etc.